# Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies

### v2

@@ Abstract
Positivity violations pose a key challenge in the estimation of causal effects, particularly for continuous interventions. Current approaches for addressing this issue include the use of weights or modified treatment policies. While effective in many contexts, these methods can result in estimands that do not always align well with the original research question, thereby compromising interpretability.
In this paper, we introduce a novel diagnostic tool—the non-overlap ratio—to detect positivity violations. To address these violations while maintaining interpretability, we propose a data-adaptive solution, specifically the “most feasible” intervention strategy.
Our strategy operates on a unit-specific basis. For a given intervention of interest, we first assess whether the intervention value is feasible for each unit. For units with sufficient support—conditional on confounders—we adhere to the intervention of interest. However, for units lacking sufficient support, we do not assign the actual intervention value of interest. Instead, we assign the closest feasible value within the support region. The non-overlap ratio provides a diagnostic summary of such support across the population.
We propose an estimator using g-computation coupled with flexible conditional density estimation to identify high- and low-support regions and to estimate this new estimand. Through simulations, we demonstrate that our method effectively reduces bias across various scenarios by addressing positivity violations. Moreover, when positivity violations are absent, the method successfully recovers the standard estimand. We further validate its practical utility using real-world data from the CHAPAS-3 trial, which enrolled HIV-positive children in Zambia and Uganda.
Keywords: data-adaptive interventions; causal inference; positivity violations; HIV treatment
@@ 1 Introduction
Causal inference methods have been extensively studied for binary treatments, where individuals are either exposed to a treatment or assigned to a control condition [21, 17]. In this framework, the positivity assumption—requiring that each individual has a non-zero probability of receiving any treatment level under consideration within confounder strata—is crucial for the identifiability and unbiased estimation of causal effects.
@@
In pharmacoepidemiological and public health studies, researchers often extend their focus to continuous interventions, where the treatment variable can take on a range of values. In this context, the estimation of the causal dose–response curve (CDRC)—which describes the relationship between the continuous treatment and the outcome—has gained increasing attention in the causal inference literature. The CDRC provides a nuanced understanding of the causal effects of continuous interventions. However, estimating the CDRC introduces new challenges, particularly because the identification of the CDRC necessitates a more restrictive positivity assumption, requiring that the conditional treatment density is greater than zero—which, when interpreted strictly, is never fully met.
For continuous interventions, the positivity assumption is commonly violated due to the infinite dimensionality of the treatment space. This makes it virtually impossible to observe all potential treatment levels across all confounder combinations. The resulting sparsity in certain regions of the intervention space can lead to unstable or biased estimates—a challenge frequently overlooked in existing studies.
@@
While methods adapted from binary treatments can be applied, they have limitations. Trimming, for example, excludes regions of the confounder space with poor overlap [7], removing individuals with extreme probabilities of treatment given their confounders. However, this approach may result in the loss of valuable data.
@@
To tackle these challenges, researchers have developed specialized methods, including projection functions and g-computation-type plug-in estimators, which aim to mitigate the violation of the positivity assumption [27]. One such approach is the weighted CDRC, which reweights and redefines the estimand of interest based on functions of the conditional support for the respective interventions [31]. Additionally, machine learning techniques have proven effective in enhancing estimation accuracy by minimizing the risk of bias caused by model misspecification [25].
@@
Another notable strategy involves modified treatment policies (MTPs)[11, 9], which specify systematic modifications to the observed (natural) treatment, with the goal of avoiding interventions among individuals for whom the treatment would be infeasible. When designed carefully, MTPs can substantially reduce reliance on the positivity assumption. However, they typically target a different causal estimand than the causal dose–response curve (CDRC), which may or may not align with the scientific objectives of a particular study—including the empirical example discussed in Section 5. MTPs are most useful when the modified intervention itself is of substantive scientific interest. Haneuse and Rotnitzky provide an example of a very well-motivated MTP [11].
@@
While double robustness and semiparametric efficiency are desirable properties, they are not always achievable when estimating the full CDRC under continuous interventions. In fact, under only mild smoothness conditions, the CDRC is not pathwise differentiable [8], precluding the construction of root-$n$ consistent or doubly robust estimators without additional structure. Recent work has developed doubly robust estimators under assumptions such as differentiability [22], monotonicity [36], or smooth invertibility [9], but these assumptions may not be justified in many applied settings—especially when the estimand is defined by the scientific question and not chosen for statistical convenience.
In this paper, we pursue a complementary goal: to preserve the original CDRC as the target estimand while improving its identifiability and estimation by redefining the intervention space in a data-adaptive way. This leads to a natural plug-in strategy that avoids extrapolation into unsupported regions, thereby enhancing both interpretability and applicability in the presence of positivity violations. Although such estimators are not doubly robust, they remain closely aligned with the scientific question, offer clear interpretability, and reduce sensitivity to model misspecification by focusing estimation in well-supported regions. Our specific contributions are:
@@
    Defining and diagnosing positivity violations in practice: we introduce a framework for assessing positivity in continuous interventions. Specifically, we define the support using the highest density region of the conditional intervention distribution and propose the non-overlap ratio—a population-level diagnostic that summarizes the extent of non-overlap across confounder strata, thereby quantifying the severity of positivity violations.
@@
    Introducing novel estimands to address positivity violations: we propose a framework of data-adaptive interventions—specifically, the most feasible intervention—to construct a new estimand that replaces infeasible interventions with the closest feasible values based on the observed data. This approach improves both interpretability and applicability, allowing researchers to approximate the standard CDRC without significantly deviating from the original intervention strategy.
By addressing the challenges posed by positivity violations in continuous interventions, our proposed methods provide a practical and reliable framework for causal inference. We demonstrate the effectiveness of our approach through simulations and real data analysis, highlighting the advantages of our method in mitigating positivity violations while preserving interpretability through minimal adjustments to the infeasible interventions.
2 From Identification Assumptions to Positivity Violations
@@ 2.1 Notation
We consider the following setup:
  • $Y$: outcome, taking values in $\mathcal{Y}$.
  • AA: treatment (or exposure), taking values in $\mathcal{A}$.
  • $\mathbf{L}=(L_{1},\dots,L_{q})$: vector of $q\in\mathbb{N}$ confounders, taking values in $\mathcal{L}$.
  • Temporal ordering: $L_{1},\dots,L_{q}\rightarrow A\rightarrow Y$.
Uppercase letters denote random variables, and lowercase letters their realizations. We adopt the potential outcome framework [29]: for subject $i$, $Y_{i}^{a}$ denotes the counterfactual outcome that would be observed if $A_{i}$ were set to level $a\in\mathcal{A}$.
Let $O=(\mathbf{L},A,Y)$ denote a generic observation, with true distribution $P_{0}$. The observed data are
$O_{1},\dots,O_{n}\;\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\;P_{0}.$
We use the following notation for distributions and densities:
  • The marginal distribution of $A$ is $P_{A}$, with density $f_{A}$.
  • The marginal distribution of $\mathbf{L}$ is $P_{\mathbf{L}}$, with density $f_{\mathbf{L}}$.
  • The conditional distribution of $A$ given $\mathbf{L}$ is $P_{A|\mathbf{L}}$, with density $f_{A|\mathbf{L}}$.
Throughout, $P_{0}$ denotes the true distribution, $P$ a generic distribution (in theoretical results), and $P_{n}$ the empirical distribution.
@@ 2.2 Determination of Positivity Violations
In causal inference, assumptions such as consistency, exchangeability, and positivity are critical for identifying causal effects. Among these, positivity—often regarded as straightforward—plays a unique role in ensuring that causal treatment effects can be properly identified.
Formally, the positivity assumption requires that every treatment level has positive probability within each stratum of the confounders $\mathbf{L}$ [27]:
$\underset{a\in\mathcal{A}}{\mathrm{inf}}\,f_{A|\mathbf{L}}(a\mid\mathbf{L})>0,\quad\text{a.e.}$ (1)
That is, the conditional treatment density must be bounded away from zero across the entire intervention space $\mathcal{A}$, with the precise requirement depending on the distribution of the intervention. While positivity is typically discussed as an abstract identification assumption, less attention has been paid to detecting violations in practice.
Even when identification holds, sparse data in certain regions of the intervention space can severely affect bias and precision at the estimation stage (see Appendix A for details). Reliable causal estimation becomes difficult when parts of the intervention domain are poorly represented for some confounder patterns, introducing bias and limiting the generalizability of the estimated causal effects.
Determining whether the data provide sufficient support for a given intervention requires assessing whether the density is so low that observing it in finite samples becomes nearly impossible. This motivates introducing a threshold $c$ on the density function to identify potential positivity violations, as proposed by Schomaker et al. [31]. Given a specific realization of confounders, $\mathbf{L}=\mathbf{l}$, an intervention of interest $a$ can be classified as:
@@
For binary interventions, (near-)violations occur when the conditional probability of observing an intervention is extremely small or zero. In this discrete case, densities coincide with probabilities, making violations relatively straightforward to detect. For continuous interventions, however, the situation is more complex. The density function must satisfy
$\int_{\mathcal{A}}f(a)\,da=1,$
so the scale of $f$ depends on the size of the intervention space $\mathcal{A}$. This variability complicates the use of static thresholds; for instance, defining interventions as unsupported when $f<0.001$ or as sufficiently supported when $f>0.1$ may not be meaningful across different contexts.
Because of this, fixed thresholds are problematic for continuous interventions. To address this limitation, we propose a novel dynamic diagnostic measure that redefines and more accurately determines positivity violations in practical settings.
@@ 2.3.1 Conditional support via highest density regions
To address the issue of poorly supported intervention areas, we propose a method that dichotomizes the normalized density values. This approach identifies regions of the intervention space $\mathcal{A}$ where support is either adequate or lacking, based on a probability threshold. Specifically, we partition $\mathcal{A}$ into two sets, $\mathcal{A}_{\alpha}(\mathbf{l})$ and its complement $\mathcal{A}^{c}_{\alpha}(\mathbf{l})$, according to a specified threshold depending on $\mathbf{l}$. Given a threshold for non-overlap detection, referred to as the support level $\alpha$, the set $\mathcal{A}_{\alpha}(\mathbf{l})$ is defined such that
$P(\mathcal{A}_{\alpha}(\mathbf{l}))=\alpha.$
An intervention $a$ can then be categorized as
[…]
We define $\mathcal{A}_{\alpha}(\mathbf{l})$ using highest density regions (HDRs), which represent the practical support of the intervention. This concept, introduced by Hyndman [20], is formally given by
$\mathcal{A}_{\alpha}(\mathbf{l})=\left\{a\in\mathcal{A}:f_{A|\mathbf{L}}(a\mid\mathbf{l})\geqslant f_{\alpha}\right\},$ (4)
where $f_{\alpha}$ is a threshold chosen such that
$P_{A|\mathbf{L}=\mathbf{l}}\!\left(\mathcal{A}_{\alpha}(\mathbf{l})\right)=\alpha.$
Figure 1 illustrates this idea: the green regions indicate the conditional support where interventions are well-supported by the data, while the grey regions highlight areas of positivity violations. In this example, the total probability of the green region corresponds to a support level of $\alpha=0.95$.
Figure 1: Illustration of using highest density regions as conditional support. Green areas indicate well-supported interventions, while grey areas correspond to positivity violations.

### v3

# Feasible Dose–Response Curves for Continuous Treatments Under Positivity Violations

@@ Abstract
Positivity violations can complicate estimation and interpretation of causal dose–response curves (CDRCs) for continuous interventions. Recently proposed weighting-based methods are designed to handle limited overlap, but the resulting weighted targets can be difficult to interpret scientifically. Modified treatment policies (MTPs) are often less sensitive to support limitations if designed properly, yet they typically target policy-defined effects that may not align with the original dose–response question. We develop an approach that addresses limited overlap while remaining as close as possible to the scientific target of the CDRC.
Our work is motivated by the CHAPAS-3 trial of HIV-positive children in Zambia and Uganda, where clinically relevant efavirenz concentration levels are not uniformly supported across covariate strata. We introduce a new diagnostic—the non-overlap ratio—which, as a function of the target intervention level, quantifies the proportion of the population for whom that level is not supported given observed covariates. We also define an individualized most feasible intervention: for each child and target concentration, we retain the target when it is supported, and otherwise map it to the nearest supported concentration. The resulting feasible dose–response curve answers: if we try to set everyone to a given concentration, but that concentration is not realistically attainable for some individuals, what outcome would be expected after shifting those individuals to their nearest attainable concentration?
We propose a plug-in g-computation estimator that combines outcome regression with flexible conditional density estimation to learn supported regions and evaluate the feasible estimand. In simulations, the method reduces bias under positivity violations and recovers the standard CDRC when support is adequate. We apply the approach to CHAPAS-3, demonstrating that feasible dose–response estimation yields a stable and interpretable concentration–response summary under realistic support constraints.
Keywords: continuous interventions; causal inference; positivity violations; HIV treatment
@@ 1 Introduction
Causal inference methods have been extensively studied for binary treatments, where individuals are either exposed to a treatment or assigned to a control condition [21, 17]. In this framework, the positivity assumption—requiring that each individual has a non-zero probability of receiving any treatment level under consideration within confounder strata—is crucial for the identifiability and estimation of causal effects.
@@
In pharmacoepidemiological and public health studies, researchers often extend their focus to continuous interventions, where the treatment variable can take on a range of values. In this context, the estimation of the causal dose–response curve (CDRC)—which describes the relationship between the continuous treatment and the outcome—has gained increasing attention in the causal inference literature. The CDRC provides a nuanced understanding of the causal effects of continuous interventions. However, estimating the CDRC introduces new challenges, particularly because the identification of the CDRC necessitates a more restrictive positivity assumption.
In this paper, “dose” is used generically to denote a continuous intervention level. Accordingly, references to the dose–response curve also encompass concentration–response and other continuous-exposure response relations.
For CDRC estimation, positivity can be more difficult to satisfy in practice because the treatment is continuous and support must be assessed over a much richer intervention space. In finite samples, some treatment regions may be sparsely represented within certain confounder strata, even when positivity is plausible at the population level. This limited overlap can reduce precision and, depending on the estimator and degree of extrapolation, may also introduce bias.
@@
While methods adapted from binary treatments can be applied, they have limitations. Trimming, for example, excludes regions of the confounder space with poor overlap [7], removing individuals with extreme probabilities of treatment given their confounders. However, this approach may result in the loss of valuable data. Moreover, it changes the target estimand: the resulting effect is interpreted conditional on the retained subset of the population, rather than the original full study population.
@@
To tackle these challenges, researchers have developed specialized methods, including projection functions, which aim to mitigate the violation of the positivity assumption [27]. One such approach is the weighted CDRC, which reweights and redefines the estimand of interest based on functions of the conditional support for the respective interventions [31]. Additionally, machine learning techniques have proven effective in enhancing estimation accuracy by minimizing the risk of bias caused by model misspecification [25].
@@
A related approach is to consider modified treatment policies (MTPs) [11, 9], which define scientifically motivated interventions by systematically transforming each individual’s observed (natural) treatment (e.g., shifting, or truncating doses). Because MTPs map observed treatment values to nearby, well-supported values, they can also improve overlap and mitigate practical positivity concerns as a secondary benefit. When designed carefully, MTPs can substantially reduce reliance on the positivity assumption. However, they typically target a different causal estimand than the causal dose–response curve (CDRC), which may or may not align with the scientific objectives of a particular study—including the empirical example discussed in Section 5. MTPs are most useful when the modified intervention itself is of substantive scientific interest. Haneuse and Rotnitzky provide an example of a very well-motivated MTP [11].
@@
While double robustness and semiparametric efficiency are often desirable, they are not generally available for estimation of the full CDRC under continuous interventions. Under mild smoothness conditions, the full CDRC functional is not pathwise differentiable [8], which precludes regular root-$n$-consistent estimation and corresponding first-order doubly robust estimators without further structure. Subsequent work has developed doubly robust estimators under additional assumptions, including differentiability of the relevant functionals [22], monotonicity restrictions [37], and smooth invertibility conditions [9]. The suitability of these assumptions is application-specific.
In this paper, we take a complementary approach. Rather than changing the target population (as in trimming) or targeting a policy parameter based on a pre-specified treatment transformation, we define a feasible dose–response curve (FDRC) using a covariate-specific feasible-assignment rule. For each attempted dose, the rule assigns that dose when it is sufficiently supported among individuals with similar covariate histories, and otherwise assigns the nearest adequately supported dose. FDRC is the mean counterfactual outcome under this covariate-adaptive intervention assignment. It therefore preserves the original dose-response relationship where overlap is adequate, while remaining identifiable in regions with limited empirical support.
Our specific contributions are:
@@
    Diagnosing practical positivity violations for continuous interventions. We develop a framework for assessing support under continuous treatment by defining covariate-specific feasible regions via highest-density regions of the conditional treatment distribution. Building on this definition, we introduce the non-overlap ratio, a population-level diagnostic that summarizes the extent of support limitations across confounder strata and quantifies the practical severity of positivity violations.
@@
    Defining the most feasible intervention and the feasible dose–response curve. We propose a feasible-intervention framework that maps infeasible target doses to nearby feasible values supported by the data. This induces a new estimand: feasible dose–response curve that remains close to the original CDRC while preserving interpretability and improving estimation under limited overlap.
Overall, the proposed framework provides a practical and scientifically coherent approach for causal inference with continuous interventions when positivity is limited. We evaluate its performance in simulations and demonstrate its utility in a real-data application, showing improved stability and reduced reliance on unsupported extrapolation while retaining a target parameter closely aligned with the original dose–response question.
The remainder of the paper is organized as follows. Section 2 presents the motivating data example. Section 3 introduces the support framework and the non-overlap ratio. Section 4 defines the proposed feasible dose–response curve (FDRC) estimand and describes its identification, estimation strategy, and implementation algorithm. Section 5 reports simulation results, and Section 6 presents the real-data application. Section 7 concludes with a discussion of implications and future directions. Section 8 illustrates the practical implementation in the CICI package.
2 Motivation Example
We motivate the problem using pharmacoepidemiological data from the CHAPAS-3 trial, which enrolled HIV-positive children in Zambia and Uganda [26, 1]. Our analysis focuses on children receiving efavirenz as part of combination antiretroviral therapy, together with lamivudine and either stavudine, zidovudine, or abacavir. Although dosing is weight-based, efavirenz concentration exposure varies substantially across children due to heterogeneity in metabolism—in particular, single nucleotide polymorphisms in the CYP2B6 gene encoding the key metabolizing enzyme—as well as adherence and other clinical factors. Because antiviral activity depends on exposure, subtherapeutic concentrations may lead to viral failure and subsequent regimen change. This motivates the scientific question: how does the probability of viral failure vary across efavirenz concentration levels?
Let EFVt denote the plasma efavirenz concentration (mg/L) at 12 hours after dosing at visit $t$, and let VLt indicate viral failure (viral load >100>100 copies/mL). A standard causal target is the concentration–response curve
$a\;\mapsto\;P\!\left(\mathrm{VL}_{t}^{\mathrm{EFV}=a}=1\right),$
defined over a clinically relevant range, e.g. $a\in(0,6)$. Informally, this estimand targets the question at which concentration levels in the blood the drug is still effective, and safe. Identification of this curve at a given level $a$ requires a positivity assumption relative to the estimand: within strata of covariates, the conditional treatment density must be bounded away from zero on the region where the curve is evaluated.
In pharmacological settings, this requirement can be difficult to satisfy because plasma concentration is not directly manipulable in the way an idealized intervention $\mathrm{EFV}_{t}=a$ suggests. Concentration is a downstream biomarker generated by a pharmacokinetic process linking dosing to absorption, distribution, metabolism, and elimination, and thus depends on individual characteristics and behaviors that are only partially controllable. For example, children with slow efavirenz clearance may be unable to achieve very low EFVt values even under clinically acceptable dose reductions, whereas fast metabolizers may rarely achieve very high concentrations at standard dosing. Weight- and age-based dosing guidelines further restrict the dose range available for a given child, and additional clinical constraints (e.g., toxicity concerns, comorbidities, drug–drug interactions, and regimen requirements) limit feasible dosing modifications. Adherence adds further structure: missed or delayed doses can yield low measured concentrations, but such values may occur primarily under specific adherence patterns and covariate profiles, making them sparse in the observed data.
When target concentration levels are weakly supported for a nontrivial subset of children, two considerations arise:
  1. 1.
    Statistical consideration. Estimation of causal effects at those levels can be sensitive to model specification and finite-sample variability, because estimation may depend more heavily on extrapolation and/or unstable weighting when estimated treatment densities are very small.
  2. 2.
    Clinical/decision consideration. Interpreting such targets as actionable interventions may be less straightforward, since some concentration levels may be difficult to achieve for many children under realistic clinical management.
These considerations motivate prioritizing intervention values with adequate empirical support and considering estimands and methods that explicitly account for practical attainability.
In our application (Section 6), empirical support for EFVt is concentrated in the mid-range (approximately 113.53.5 mg/L), with substantially weaker support near 0 mg/L and above about $3.5$ mg/L, indicating limited overlap for target intervention values at the extremes (Figure 6a). Estimation of $P(\mathrm{VL}_{t}^{\mathrm{EFV}=a}=1)$ at such extremes may therefore rely more heavily on extrapolation, and the estimated feasible dose-response curve departs from the standard concentration–response curve primarily in these regions, while remaining similar where support is adequate (Figure 6b6c).
3 Positivity and Diagnostics
@@ 3.1 Notation
Let $O=(\mathbf{L},A,Y)\in\mathcal{O}:=\mathcal{L}\times\mathcal{A}\times\mathcal{Y}$ denote one observed data unit, where
  • $Y\in\mathcal{Y}\subseteq\mathbb{R}$ is the outcome,
  • $A\in\mathcal{A}\subseteq\mathbb{R}$ is the continuous treatment (or exposure),
  • $\mathbf{L}=(L_{1},\dots,L_{q})\in\mathcal{L}\subseteq\mathbb{R}^{q}$ is a vector of baseline covariates sufficient to control confounding.
We assume the temporal ordering
$L_{1}\rightarrow L_{2}\rightarrow\cdots\rightarrow L_{q}\rightarrow A\rightarrow Y,$
where each $L_{j}$ is measured pre-treatment.
The observed sample is
$O_{1},\dots,O_{n}\overset{\mathrm{iid}}{\sim}P_{0},$
where $P_{0}$ is the true data-generating distribution. Uppercase letters denote random variables and lowercase letters their realizations.
For each $a\in\mathcal{A}$, let $Y^{a}$ be the potential outcome [29] under the intervention that sets treatment to $a$. The causal dose–response curve is
$m_{0}(a):=\mathbb{E}_{0}\!\left[Y^{a}\right],\qquad a\in\mathcal{A}.$ (1)
When densities exist (with respect to appropriate dominating measures), we use:
  • $P_{0,A}$ and density $f_{0,A}$ for the marginal law of $A$,
  • $P_{0,\mathbf{L}}$ and density $f_{0,\mathbf{L}}$ for the marginal law of $\mathbf{L}$,
  • $P_{0,A\mid\mathbf{L}=\mathbf{l}}$ and density $f_{0,A\mid\mathbf{L}}(a\mid\mathbf{l})$ for the conditional law of $A\mid\mathbf{L}=\mathbf{l}$.
Here and throughout, $P$ denotes a generic distribution in the statistical model $\mathcal{M}$, and $P_{n}$ denotes the empirical measure.
@@ 3.2 Determination of Positivity Violations
Identification of the causal dose–response curve in (16) relies on the standard assumptions of consistency, exchangeability, and positivity. A common positivity condition [27] is
$\inf_{a\in\mathcal{A}}f_{0,A\mid\mathbf{L}}(a\mid\mathbf{L})>0,\quad\text{a.e.}$ (2)
where “a.e.” is with respect to $P_{0,\mathbf{L}}$. Equivalently, for $P_{0,\mathbf{L}}$-almost every $\mathbf{l}$, all treatment levels $a\in\mathcal{A}$ relevant to (16) have strictly positive conditional density under $A\mid\mathbf{L}=\mathbf{l}$.
While (2) is a population-level condition, empirical estimation depends on finite-sample support. In practice, some intervention levels may be weakly represented within certain covariate strata, which can reduce precision and increase reliance on extrapolation (see Appendix A). Thus, a practical diagnostic of support is needed.
A natural starting point is to classify support using a density threshold $c>0$, as considered in Schomaker et al. [31]. For fixed $(a,\mathbf{l})$, define
@@
This formalizes the idea that near-zero conditional density corresponds to weak support.
For binary treatments, this assessment is comparatively straightforward because conditional densities coincide with conditional probabilities. For continuous treatments, however, fixed cutoffs are less portable across settings: for each $\mathbf{l}$,
$\int_{\mathcal{A}}f_{0,A\mid\mathbf{L}}(a\mid\mathbf{l})\,da=1,$
so the numerical scale of the density depends on the measurement scale and effective range of $\mathcal{A}$. Consequently, universal thresholds (e.g., $f_{0,A\mid\mathbf{L}}(a\mid\mathbf{l})<0.001$) are difficult to interpret consistently across studies.
Motivated by this limitation, we next introduce a diagnostic based on covariate-specific highest-density regions, providing a scale-aware characterization of practical positivity violations for continuous interventions.
@@ 3.3.1 Conditional support via highest density regions
To characterize practical support for intervention values conditional on covariates, fix support level $\alpha\in(0,1)$ and define a covariate-specific set $\mathcal{A}_{\alpha}(\mathbf{l})\subseteq\mathcal{A}$. Intuitively, $\mathcal{A}_{\alpha}(\mathbf{l})$ contains treatment values with relatively high conditional density under $A\mid\mathbf{L}=\mathbf{l}$, chosen to contain $\alpha$ of the conditional probability mass. We refer to $a\in\mathcal{A}_{\alpha}(\mathbf{l})$ as well supported at level $\alpha$; when no ambiguity arises, we write simply supported. Values in $\mathcal{A}\setminus\mathcal{A}_{\alpha}(\mathbf{l})$ are poorly supported (or unsupported).
Formally, let $f_{0,A\mid\mathbf{L}}(a\mid\mathbf{l})$ be a version of the conditional density of $A\mid\mathbf{L}=\mathbf{l}$ (with respect to a dominating measure on $\mathcal{A}$). Following the HDR construction of Hyndman [20], define a cutoff $f_{0,\alpha}(\mathbf{l})$ (Appendix C) and set
$\mathcal{A}_{\alpha}(\mathbf{l})=\{a\in\mathcal{A}:\ f_{0,A\mid\mathbf{L}}(a\mid\mathbf{l})\geq f_{0,\alpha}(\mathbf{l})\}$ (4)
such that
$P_{0,A\mid\mathbf{L}=\mathbf{l}}\!\big(\mathcal{A}_{\alpha}(\mathbf{l})\big)=\alpha.$
Thus, for fixed $\alpha$, intervention values outside $\mathcal{A}_{\alpha}(\mathbf{l})$ correspond to relatively low conditional density and are most prone to practical positivity violations.
If $f_{0,A\mid\mathbf{L}}(a\mid\mathbf{l})$ is constant on $\mathcal{A}$ (e.g., $A\mid\mathbf{L}=\mathbf{l}$ is uniform), the HDR is non-unique: any measurable subset of $\mathcal{A}$ with conditional probability $\alpha$ satisfies (4). In that case, the well-supported/poorly-supported distinction is not informative.
Figure 1 illustrates the construction for $\alpha=0.95$: the green shaded area is the conditional probability mass inside $\mathcal{A}_{0.95}(\mathbf{l})$, and the grey area is the remaining mass outside.
Figure 1: Illustration of the highest density regions (HDRs) used to define covariate-specific intervention support. For fixed $\mathbf{l}$ and $\alpha=0.95$, well-supported intervention values are those on the $x$-axis in $\mathcal{A}_{\alpha}(\mathbf{l})$. The green shaded region indicates probability mass of $A\mid\mathbf{L}=\mathbf{l}$ inside $\mathcal{A}_{\alpha}(\mathbf{l})$; the grey region indicates mass outside.